### v1

# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

Abstract
The superconducting diode effect (SDE) refers to the non-reciprocal nature of the critical current (maximum current that a superconductor can withstand before turning into a normal metal) of a superconducting device. Here, we investigate SDE in helical superconductors with broken inversion and time-reversal symmetry, focusing on a prototypical Rashba nanowire device proximitized by an $s$-wave superconductor and subjected to external magnetic fields. Using a self-consistent Bogoliubov-de Gennes mean-field formalism, we analyze the interplay between linear and higher-order spin-orbit coupling (SOC), bulk supercurrents, and external magnetic fields. Our results demonstrate that Rashba nanowires with only linear SOC can achieve incredibly large diode efficiencies of $\gtrsim 45\%$ through the interplay of longitudinal and transverse magnetic fields. Notably, higher-order SOC introduces qualitatively different behavior, enabling finite diode efficiency even in the absence of a longitudinal Zeeman field due to inherent energy dispersion asymmetry. We present a comprehensive phase diagram of the device elucidating the emergent Fulde-Ferrell-Larkin-Ovchinnikov (FFLO) superconducting state and demonstrate that proximitized Rashba nanowires offer a versatile, practical platform for SDE, with potential realizations in existing material systems. These results provide crucial insights for optimizing SDE in nanoscale superconducting devices, paving the way for next-generation dissipationless quantum electronics.
I Introduction
The invention of diodes marked a pivotal breakthrough in solid-state electronics, with their distinctive non-reciprocal current flow making them fundamental to semiconductor devices Scaff and Ohl (1947); Shockley (1949). In quantum materials, inversion and time-reversal symmetry crucially determine charge current nonreciprocity. Magnetochiral anisotropy (MCA), arising in systems with broken inversion symmetry Tokura and Nagaosa (2018), is a dominant mechanism for this effect. MCA leads to non-reciprocity in the magnetoresistance with respect to the current direction and can be greatly enhanced by superconducting correlations Qin et al. (2017); Wakatsuki et al. (2017). In superconductors, MCA can lead to an extreme manifestation of nonreciprocity known as the superconducting diode effect (SDE), where resistance is zero in one current direction and non-zero in the other Nadeem et al. (2023); Nagaosa and Yanase (2024); Narita and Ono (2024). SDE thus realizes an ideal diode with perfect rectification. Such a scenario arises in the regime where the critical current that causes superconductor to normal metal transition differ in the opposite directions ($|J_{c}^{+}|\neq|J_{c}^{-}|$Nadeem et al. (2023); Nagaosa and Yanase (2024). Recent experiments in engineered superlattice systems Ando et al. (2020); Sundaresh et al. (2023); Narita and Ono (2024) and bulk materials/ thin films Itahashi et al. (2020); Schumann et al. (2020); Wakatsuki et al. (2017); Nadeem et al. (2023) have spurred research interests into more efficient superconducting diode devices, including in twisted multilayer graphene Lin et al. (2022); Díez-Mérida et al. (2023) and transition metal dichalcogenides Bauriedl et al. (2022); Yun et al. (2023).
The inversion symmetry breaking required for critical current non-reciprocity can be achieved through extrinsic mechanisms (e.g., asymmetric geometry or artificial superlattice potential Sundaresh et al. (2023); Burlakov et al. (2014); Narita and Ono (2024)) or intrinsic mechanisms like Rashba spin-orbit coupling (SOC) Ando et al. (2020); Yuan and Fu (2022); Ilić and Bergeret (2022); Halterman et al. (2022); Daido and Yanase (2022); Karabassov et al. (2022); Legg et al. (2022). Correspondingly, SDE can be broadly classified into two types: extrinsic and intrinsic. Although predicted decades ago Levitov et al. (1985); Edelstein (1996), the mechanism of SDE has only recently been clearly understood, leading to extensive investigations into its potential for dissipationless electronic circuits Yuan and Fu (2022); Daido et al. (2022); Ilić and Bergeret (2022); He et al. (2022); Pal et al. (2022); de Picoli et al. (2023); Banerjee and Scheurer (2024); Nadeem et al. (2023). Recent theoretical studies Yuan and Fu (2022); Daido et al. (2022); Ilić and Bergeret (2022); He et al. (2022), using phenomenological Ginzburg-Landau theory and mean-field analysis of the two-dimensional Rashba-Zeeman Hubbard model, have shown that intrinsic SDE can arise in systems with broken inversion and time-reversal symmetry. In the helical superconducting ground state, this occurs due to asymmetry in the FFLO Cooper pair momentum, resulting in Kramers non-degenerate $s$-wave Cooper pairs Daido et al. (2022); Yuan and Fu (2022); Nadeem et al. (2023).
A Rashba nanowire device proximitized to an s-wave superconductor and subjected to external magnetic fields serves as a prototypical platform for realizing helical superconductivity Oreg et al. (2010); Sau et al. (2010); Lutchyn et al. (2010); Stanescu et al. (2011); Legg et al. (2022); de Picoli et al. (2023). While previous phenomenological/ non-self-consistent studies Legg et al. (2022) demonstrated the potential for superconducting diode effect (SDE) in such systems, they reported low diode efficiencies ($\sim 2\%$). However, to accurately determine the diode efficiency in the true superconducting ground state across the parameter space, it is crucial to systematically investigate the superconducting properties self-consistently, accounting for variations in the superconducting order parameter, as demonstrated in recent studies Daido et al. (2022); Ilić and Bergeret (2022); de Picoli et al. (2023). In this paper, we address this gap through a comprehensive, self-consistent mean-field investigation of the interplay between various Rashba SOC terms and external magnetic fields in determining emergent FFLO pairing and highly efficient superconducting diodes in a Rashba nanowire device. We show that even with only linear Rashba SOC we can achieve a very large diode efficiency ($\gtrsim 45\%$) by suitable tuning the parameters in this system. We extend beyond linear Rashba SOC to include higher-order (cubic) Rashba SOC terms, which are not only naturally present in semiconducting nanowires Campos et al. (2018); Kammermeier et al. (2018); Sonehara et al. (2022) but also can dominate the normal state properties Vajna et al. (2012); Moriya et al. (2014); Nakamura et al. (2012); Michiardi et al. (2015); Shanavas (2016); Liu et al. (2018); Peng et al. (2023); Amundsen et al. (2024). Notably, we demonstrate that higher-order Rashba SOC can induce finite diode efficiency ($\gtrsim 15\%$) even without longitudinal magnetic fields. Our study establishes the Rashba nanowire device as a promising platform for highly efficient superconducting diodes, crucial for advancing nanoscale superconducting electronics.
The remainder of this article is structured as follows. Sec. II introduces the proximity-induced Rashba nanowire device and outlines the self-consistent mean-field formalism used to investigate properties of the superconducting ground state in the nanowire. Sec. III presents our main findings on nonreciprocal charge transport in this device. We examine cases with both linear and higher-order Rashba SOC terms, exploring their interplay with Zeeman fields to optimize the SDE. Finally, Sec. IV summarizes our results, discusses potential material realizations, and proposes future research directions.
Figure 1: Schematic of the proximity-induced Rashba nanowire device: (a) The schematic of our proposed device consisting of a Rashba nanowire in close proximity to an $s$-wave superconductor and in presence of external Zeeman fields: $B_{x}$ and $B_{z}$ to investigate SDE. The strength of the linear Rashba SOC is denoted by $\alpha$, while the cubic Rashba SOC strengths are represented by $\delta$ and $\gamma$ as illustrated in Eq. (2). (b) and (c): The dispersions of the normal state and the Bogoliubov quasi particles in the superconducting state are qualitatively shown in (b) and (c) respectively considering both linear and higher order SOC.
II Model and Self-consistent mean-field formalism
We consider a single channel one-dimensional semiconducting nanowire with both linear and higher order Rashba SOC Campos et al. (2018); Kammermeier et al. (2018); Sonehara et al. (2022) and in presence of externally applied magnetic fields leading to the Rashba-Zeeman nanowire Hamiltonian (Alicea, 2012)
[…]
$\!\!\hat{h}(z)=\frac{p_{z}^{2}}{2m}-\mu+\left(\frac{\alpha}{\hbar}p_{z}+\frac{\delta}{\hbar^{3}}p_{z}^{3}\right)\sigma_{x}+\frac{\gamma}{\hbar^{3}}p_{z}^{3}\sigma_{y}+\mathbf{B}.\boldsymbol{\sigma}$ (1)
Here, the spin degree of freedom is labeled by $s,s^{\prime}\in\{\uparrow,\downarrow\}$. The nanowire is placed along the $z$-direction (as shown in Fig. 1(a)) and $\psi_{s}^{\dagger}(z)$ operator creates a fermion at the position $z$ with spin $s$. $\boldsymbol{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is the vector of the Pauli matrices in the spin space and $\mu$ is the chemical potential. $p_{z}=\frac{\hbar}{i}\partial_{z}$ is the momentum operator in the direction of current flow ($z$-direction) within the nanowire. The strength of the usual linear Rashba SOC is denoted by $\alpha$ and the cubic Rashba SOC of strengths are $\delta$ and $\gamma$. The presence of Rashba SOC breaks the inversion symmetry of the system. The externally applied magnetic fields lead to the Zeeman term characterized by $\mathbf{B}=(B_{x},B_{y},B_{z})$ that breaks time-reversal symmetry. We assume the radius of the nanowire to be small such that there is only one occupied mode, i.e.,  single channel. This nanowire Hamiltonian [Eq. (1)] in momentum space can be written as
$\mathcal{H}_{0}=\sum_{s,s^{\prime}}\sum_{k}\,c_{k,s}^{\dagger}h_{ss^{\prime}}(k)c_{k,s^{\prime}},$
$\hat{h}(k)=\xi_{k}+\left[\left(\alpha k+\delta k^{3}\right)\sigma_{x}+\gamma k^{3}\sigma_{y}\right]+\mathbf{B}.\boldsymbol{\sigma}$ (2)
Here, $k$ represents the momentum along the nanowire ($z$-direction) and $\xi_{k}=\frac{\hbar^{2}k^{2}}{2m}-\mu$. The bulk energy spectrum of the nanowire obtained by diagonalizing $\hat{h}_{k}$ is given by
$\epsilon_{\pm}(k)=\xi_{k}\pm\sqrt{(\alpha k+\delta k^{3}+B_{x})^{2}+(\gamma k^{3}+B_{y})^{2}+B_{z}^{2}}$
where, $\pm$ label the two helicty bands. Note that, the Zeeman field $B_{z}$ opens a Zeeman gap while $B_{x}$ creates asymmetry between the two bands as schematically shown in Fig. 1(b). On the other hand, $B_{y}$ opens up a gap as well as creates band asymmetry (not shown). Throughout the majority of this paper, we will discuss the results with $B_{y}=0$ for simplicity and will separately discuss the effects of nonzero $B_{y}$ for completeness.
We seek to elucidate the nature of the superconducting ground state in the Rashba nanowire in the presence of $s$-wave superconducting correlations that can be induced via the proximity effect in presence of a bulk $s$-wave superconductor, as well as supercurrents. Such a device is schematically shown in Fig. 1(a). In general, the presence of supercurrents can lead to a superconducting state with finite momentum Cooper pairs termed as the FFLO state. To establish that, we consider an attractive Hubbard type interaction in the bulk $s$-wave superconductor described by the Hamiltonian
$\mathcal{H}_{I}=-\frac{U}{2}\sum_{s,s^{\prime}}\int d^{3}{\bf{r}}\,\psi_{s}^{\dagger}({\bf{r}})\psi_{s^{\prime}}^{\dagger}({\bf{r}})\psi_{s^{\prime}}({\bf{r}})\psi_{s}({\bf{r}})\ ,$ (3)
where, $U$ is the attractive Hubbard interaction strength. We now decouple the Hubbard interaction term ($\mathcal{H}_{I}$) in the finite momentum $s$-wave channel within the mean-field approximation Daido et al. (2022); Daido and Yanase (2022); de Picoli et al. (2023) and consider these correlations to be proximity induced in the nanowire. Then, using the Bogoliubov de-Gennes (BdG) mean-field formalism the effective Hamiltonian of the nanowire takes the form
[…]
$\!\!\!\!\!\!\!\!\!\!\mathcal{H}_{BdG}(z)$ $=$ $\begin{bmatrix}\hat{h}(z)&\hat{\Delta}(z)\\ -\hat{\Delta}^{*}(z)&-\hat{h}^{*}(z)\end{bmatrix};\hat{\Delta}(z)=-i\sigma_{y}\Delta e^{iqz},$ (4)
where, $\Psi(z)=[\psi_{\uparrow}(z),\psi_{\downarrow}(z),\psi_{\uparrow}^{\dagger}(z),\psi_{\downarrow}^{\dagger}(z)]^{T}$ is the Nambu spinor inside the nanowire, $\mathcal{E}_{0}=\frac{L}{U}|\Delta|^{2}$ with $L$ denoting the length of the nanowire. The $s$-wave FFLO order parameter is given by $\Delta(z)=\Delta(-i\sigma_{y})e^{iqz}$, where $q$ is the Cooper pair momentum. The Hamiltonian in Eq. (4) effectively describes the physics of the Rashba nanowire device schematically shown in Fig. 1(a). To proceed further, we write the mean-field Hamiltonian [Eq. (4)] in the momentum space as
$\mathcal{H}$ $=$ $\frac{1}{2}\sum_{k}\Psi_{k}^{\dagger}\mathcal{H}_{BdG}(k)\Psi_{k}+\mathcal{E}_{0},$
$=$ $\sum_{n,k}\left(E_{n,k}\gamma^{\dagger}_{n,k}\gamma_{n,k}-\frac{E_{n,k}}{2}\right)+\mathcal{E}_{0},$
[…]
Here, $\Psi_{\mathbf{k}}\equiv[c_{k+q/2,\uparrow},c_{k+q/2,\downarrow},c_{-k+q/2,\uparrow}^{\dagger},c_{-k+q/2,\downarrow}^{\dagger}]^{T}$ and $E_{n,k}(q)$ denotes the Bololiubov quasiparticle energies. The behavior of the Bogoliubov quasiparticle spectrum is schematically shown in Fig. 1(c). This depicts that the introduction of the pairing potential $\Delta$ opens a gap in the spectrum. As in the case of the nanowire spectrum, the asymmetry in BdG bands occurs due to the application of the Zeeman field $B_{x}$, and the Zeeman field $B_{z}$ makes the bands distinctly gapped.
Figure 2: FFLO ground state with linear SOC: (a) The self-consistent superconducting gap $\Delta(q)$ is shown as a function of the Cooper-pair momentum $q~{}$(nm-1) for $B_{x}/\Delta_{0}=0.3$ with only linear Rashba SOC ($\alpha=1$ eV-Å). $\Delta_{0}$ is the BCS gap in the absence of any external magnetic field ($\mathbf{B}=0$). (b) The Cooper pair momentum of the FFLO ground state $q_{0}$ as a function of $B_{x}$. The presence of the FFLO ground state ($q_{0}\neq 0$) is established only when both $B_{x}$ and $B_{z}$ are finite. Other model parameters are chosen as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $25.55$ meV, $0.1$ meV).
Figure 3: FFLO ground state with both linear and higher order SOC: (a) The self-consistent superconducting gap $\Delta(q)$ is depicted as a function of Cooper-pair momentum $q~{}$ for $B_{z}=0$ with both the linear Rashba SOC term ($\alpha=1$ eV-Å) and the cubic Rashba SOC terms ($\delta=\gamma=100$ eV-Å3). (b) The Cooper pair momentum of the FFLO ground state $q_{0}~{}$ as a function of $B_{x}$ is shown. Note that the FFLO state is stable even when $B_{z}=0$ in the presence of a finite $B_{x}$. Other system parameters are chosen as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $16.45$ meV, $0.1$ meV).
The condensation energy $\Omega(q)$ of the system, defined as the difference of free energy per unit length between the superconducting and normal state Kinnunen et al. (2018); Daido et al. (2022), is
$\Omega(q,\Delta)=F(q,\Delta)-F(q,0),$ (6)
where, $F(q,\Delta)$ is the free energy density of the nanowire given by
$F(q,\Delta)=-\frac{1}{L\beta}\sum_{n,k}\ln\left[1+e^{-\beta E_{n,k}(q)}\right]+\frac{|\Delta(q)|^{2}}{U},$ (7)
where, $\beta=(k_{B}T)^{-1}$ with $k_{B}$ being the Boltzmann constant and $T$ is the temperature.
Figure 4: Superconducting diode effect with linear SOC: (a) The supercurrent density ($J$) is shown as a function $q~{}$(nm-1) considering solely linear Rashba SOC and $B_{x}/\Delta_{0}=0.3$. The nonreciprocity in the supercurrent, $J(q)\neq-J(-q)$, is achieved in this case only when both $B_{x}$ and $B_{z}$ are finite. It then leads to nonreciprocity in the critical supercurrents, $J_{c}^{+}\neq|J_{c}^{-}|$, hence exhibiting a finite diode efficiency ($\eta\neq 0$). Here, $J_{0}\equiv J_{c}(B_{z}=0)=|J_{c}^{+}(B_{z}=0)|=|J_{c}^{-}(B_{z}=0)|$. (b) and (c): The corresponding diode efficiency ($\eta$) is shown as a function of $B_{x}$ (with fixed $B_{z}$) and $B_{z}$ (with fixed $B_{x}$) with only linear Rashba SOC in the panels (b) and (c) respectively. The system parameters are chosen as: $(B_{y},\alpha,\mu/\Delta_{0},U,\beta^{-1})$=(00, $1$eV-Å, $0.5$, $25.55$ meV, $0.1$ meV).
We determine the order parameter $\Delta(q)$ for a given value of $q$ by self-consistently solving the gap equation, obtained via minimizing the free energy $F(q,\Delta)$ in Eq. (7), given by
$\Delta(q)=-\frac{U}{L}\sum_{n,k}\frac{\partial E_{n,k}}{\partial\Delta^{*}}\,n_{F}(E_{n,k})$ (8)
where $n_{F}(E_{n,k})=1/(1+e^{\beta E_{n,k}})$ is the Fermi function. Using the self-consistent solution of $\Delta(q)$ in Eq. (6), we optimize the condensation energy with respect to $q$ to obtain the Cooper pair momentum ($q_{0}$) of the true FFLO superconducting ground state. The BCS gap in the absence of any external magnetic fields ($\mathbf{B}=0$) is denoted by $\Delta_{0}$ and we choose the value of $U$ in such a way that we are in the weak coupling BCS regime with $\Delta_{0}\sim 1$ meV. In the rest of the paper, we present results in the low temperature limit choosing a fixed value of $\beta^{-1}=0.1$ meV much smaller than $\Delta_{0}$.
The variation of the self-consistent solutions of the superconducting order parameter $\Delta(q)$ and the momentum of the Cooper pairs $q_{0}$ in the FFLO ground state with the external Zeeman fields are shown in Fig. 2 and Fig. 3 with only linear SOC case and, with both the linear and higher order SOC case respectively. We note from Fig 2(a) and Fig 3(a) that for both the cases the superconducting gap vanishes for a critical value of the Cooper pair momentum either above a positive value of $q=q_{c}^{+}$ or below a negative value of $q=q_{c}^{-}$. Fig 2(b) and Fig. 3(b) show that $q_{0}$ varies linearly with $B_{x}$ for $B_{x}\ll\Delta_{0}$ for a given value of $B_{z}$. We note Fig 2(b) that in the case of only linear SOC $q_{0}$ is finite only when $B_{z}\neq 0$ and $B_{x}\neq 0$. This is because when only linear SOC is present the effect of the SOC can be gauged away by using a spin-dependent gauge transformation Braunecker et al. (2010) and then when only $B_{x}\neq 0$ the band structure is symmetric leading to zero momentum Cooper pairs. However, when both $B_{x}$, $B_{z}\neq 0$, then in the transformed basis a spiral magnetic field texture is realized leading to momentum dependent asymmetry in the band structure and hence leading to finite momentum Cooper pairs. In contrast, when both the linear and cubic SOC terms are present, the effect of SOC can not be gauged away and as a result only $B_{x}\neq 0$ is sufficient to give rise to $q_{0}\neq 0$ in this case. Consequently, this feature also gets manifested in realizing finite SDE in this Rashba nanowire device, as extensively discussed in the next section.
Figure 5: Superconducting diode effect with both linear and higher order SOC: (a) The supercurrent density ($J$) as a function of $q~{}$(nm-1) with both cubic Rashba SOC ($\delta$ = $\gamma$ = $100$ eV-Å3) and linear Rashba SOC ($\alpha=1$ eV-Å) choosing $B_{z}/\Delta_{0}=0$. The nonreciprocity in the critical supercurrents: $|J_{c}^{+}|\neq|J_{c}^{-}|$ (also implying a finite diode efficiency $\eta\neq 0$) is achieved for any finite value of $B_{x}$. Here, $J^{\prime}_{0}\equiv J_{c}(B_{x}=B_{z}=0)=|J_{c}^{+}(B_{x}=B_{z}=0)|=|J_{c}^{-}(B_{x}=B_{z}=0)|$. (b) and (c): The corresponding diode efficiency ($\eta$) is depicted as a function of $B_{x}$ (with fixed $B_{z}$) and $B_{z}$ (with fixed $B_{x}$) in the panels (b) and (c) respectively. We have chosen the other model parameters as: $(B_{y},\mu/\Delta_{0},U,\beta^{-1})$=(00, $0.5$, $16.45$ meV, $0.1$ meV).
III Non-reciprocal transport
The supercurrent density $J(q)$ is computed from the condensation energy $\Omega(q,\Delta)$ in Eq. (6) by introducing a vector potential $A$ using minimal coupling and then taking the derivative of the condensation energy with respect to $A$ Yuan and Fu (2022) as,
$J(q)=-\frac{\partial\Omega(q+2eA,\Delta)}{\partial A}\bigg{|}_{A=0}=-2e\,\frac{\partial\Omega(q,\Delta)}{\partial q}.$ (9)
The sign of the $J(q)$ indicates the direction of the supercurrent flow. The critical values of the supercurrent along and opposite to the flow direction within the range $q_{c}^{-}\leq q\leq q_{c}^{+}$ for which the superconducting phase is stable, are denoted as $J_{c}^{+}$ and $J_{c}^{-}$ respectively. When $|J_{c}^{+}|\neq|J_{c}^{-}|$, we have a finite SDE in the system and the efficiency or the quality factor of the SDE is defined as
$\eta=\frac{|J_{c}^{+}|-|J_{c}^{-}|}{|J_{c}^{+}|+|J_{c}^{-}|}\ .$ (10)
We evaluate the diode efficiency ($\eta$) numerically, by first computing $J(q)$ over the range $q_{c}^{-}\leq q\leq q_{c}^{+}$ using the the self-consistency solutions $\Delta(q)$ for each value of $q$ shown in Fig 2(a) and Fig 3(a) for example, and then finding its optimal values $J_{c}^{+}$ and $J_{c}^{-}$. We systematically analyze the SDE in our Rashba nanowire device considering two broadly distinct scenarios: with only linear SOC and, with both linear and higher order SOC for various system parameters as discussed in the following two subsections.
III.1 Case-I: Rashba nanowire with only linear SOC
The behavior of the supercurrent density $J(q)$ and the efficiency ($\eta$) of the SDE in the case when only linear Rashba SOC is present for different system parameters are shown in Fig. 4. As shown in Fig. 4(a), the supercurrent density is symmetric with respect to $q$ for $B_{z}=0$ even with $B_{x}\neq 0$ Legg et al. (2022). We note that as we start increasing the magnitude of $B_{z}$ with $B_{x}\neq 0$, the $J(q)$ becomes more and more asymmetric, leading to nonreciprocal critical supercurrents, $|J_{c}^{+}|\neq|J_{c}^{-}|$ and hence to finite SDE. This is a direct consequence of the fact that the presence of FFLO state and the asymmetry in critical momenta: $|q_{c}^{+}|\neq|q_{c}^{-}|$ is distinctly realized only in presence of both the fields $B_{x},B_{z}$ as noted in Fig. 2(a).
The diode efficiency ($\eta$) as a function of $B_{x}$ for fixed values of $B_{z}$ and as a function of $B_{z}$ for fixed values of $B_{x}$ are shown in Fig. 4(b) and Fig. 4(c) respectively. In both cases, $\eta$ increases linearly at low field strengths, reaches a maximum at moderate values, and then decreases to zero as the fields intensify further. This behavior can be understood as a consequence of a competition between the asymmetry in the band structure of the system and changes in the superconducting pairing susceptibility as the magnetic field strengths are varied, consistent with previous studies Yuan and Fu (2022). Note that, the self-consistent BCS gap vanishes for very large values of the Zeeman fields $B_{x}$ and $B_{z}$ consequently leading to zero efficiency, however the resultant nanowire spectrum might still be gapped like an ordinary insulator.
Preceding investigations on a Rashba nanowire featuring only linear Rashba SOC Legg et al. (2022) reported maximum diode efficiency of $\sim 2\%$, without taking into consideration self-consistency in the superconducting order parameter and fixing it to a particular value. In our study, we have performed an extensive analysis of the full parameter space self-consistently and demonstrated that the diode efficiency in the true superconducting ground state of the device can be increased to $\eta\gtrsim 45\%$ [see Fig. 4(b, c)]. The maximum values of the Zeeman fields ($B_{x},B_{z}$) were restricted to the point where the self-consistent BCS gap vanishes. We note that in this case $B_{y}$ also plays a similar role as $B_{z}$ due to the geometry of the setup (Fig. 1(a)) under consideration and hence further analysis considering $B_{y}\neq 0$ is not presented. However, $B_{y}$ plays a significant and distinct role when we consider the effect of cubic Rashba SOC terms as discussed in the next subsection.
Figure 6: Diode efficiency phase diagram with both linear and higher order SOC: The variation of the diode efficiency $\eta$ is shown in the plane of cubic Rashba SOC parameters $\delta-\gamma$ choosing “optimal” values of the Zeeman fields $(B_{x}/\Delta_{0},B_{z}/\Delta_{0})=(0.4,0.0)$ in panel (a) and $(B_{x}/\Delta_{0},B_{z}/\Delta_{0})=(0.4,0.5)$ in panel (b), respectively. We note a small enhancement in the maximum value of $\eta$ for finite $B_{z}$. The other system parameters are chosen as: $(B_{y},\alpha,\mu/\Delta_{0},\beta^{-1})$=(0(0( 0, $1$ eV-Å, $0.5$, $0.1$ meV).
III.2 Case-II: Rashba nanowire with both linear and higher order SOC
The variation of the supercurrent density $J(q)$ and the diode efficiency ($\eta$) in the case when both the linear ($\alpha\neq 0$) and cubic Rashba SOC terms ($\delta,\gamma\neq 0$) are present for different system parameters are shown in Fig. 5. From the variation of $J(q)$ shown in Fig. 5(a), we note that in this case even with $B_{z}=0$ as long as $B_{x}\neq 0$ the supercurrent becomes asymmetric leading to nonreciprocal critical supercurrents, $|J_{c}^{+}|\neq|J_{c}^{-}|$ and hence to finite SDE. The variation of $\eta$ with the magnetic fields are shown in Fig. 5(b, c). We note from Fig. 5(b) that the diode efficiency $\eta$ at a fixed $B_{z}$ exhibits a sharp linear increase as a function of $B_{x}$ for small values of $B_{x}\ll\Delta$, eventually attaining a maxima around $B_{x}/\Delta_{0}=0.4$ similar to the only linear SOC case described earlier. The maximum efficiency realized increases as one increases the magnitude of $B_{z}$. However, $\eta$ shows only marginal enhancement at a fixed value of $B_{x}$ when $B_{z}$ is varied as depicted in Fig. 5(c).
Figure 7: (a) The diode efficiency $\eta$ is shown as a function of $B_{y}$ (with fixed $B_{x}/\Delta_{0}=0.4$) for our system in consideration with cubic Rashba SOC: $(\delta,\gamma)=100$ eV-Å3. (b) The $\eta$ as a function of $B_{x}$ is shown with linear SOC and with only one cubic Rashba SOC term $\delta=100$ eV-Å3 but $\gamma=0$ for different values of $B_{z}$ taking $B_{y}=0$. We have chosen the model parameters as: $(\alpha,\mu/\Delta_{0},\beta^{-1})$=($1$ eV-Å, $0.5$, $0.1$ meV).
Modifying the values of $\delta$ and $\gamma$ evidently changes the band dispersions of the system and consequently the efficiency of the superconducting diode will also change. We depict the diode efficiency phase diagrams in the $\delta-\gamma$ plane in Fig. 6. We obtain the phase diagrams by varying $U$ for every configuration of ($\delta,\gamma$) such that the self-consistent BCS gap $\Delta_{0}$ remains fixed. For the $B_{z}=0$ and $B_{x}\neq 0$ case, we can achieve a peak diode efficiency of approximately $13.6\%$ for an optimal choice of higher-order Rashba SOC parameters as shown in Fig. 6(a). Furthermore, the application of an additional $B_{z}$ field readily elevates $\eta$ to approximately $19.6\%$, as illustrated in Fig. 6(b).
Note that the results discussed so far were obtained by keeping $B_{y}=0$. Now, we would like to consider the effect of finite $B_{y}$ as well. In this case, $B_{y}$ serves the role of both $B_{x}$ and $B_{z}$ in realizing a finite SDE in our system by creating both gap and asymmetry in the band structure. We note that the application of $B_{y}$ produces linear increase in $\eta$ with increasing $B_{y}$ for fixed $B_{x}$ and $B_{z}$ as shown in the Fig. 7(a).
Finally, as a special case, we consider the case when $\gamma=0$ but $\alpha,\delta\neq 0$. We can still obtain a finite diode efficiency with a maximum of $12\%$ purely in presence of a finite $B_{x}$ even when $B_{z}=0$ [see Fig. 7(b)]. The inclusion of a single cubic Rashba SOC parameter $\delta$, in a system featuring linear Rashba SOC ($\alpha$), is sufficient to manifest SDE with exclusively finite $B_{x}$ (in contrast to the case where the system features only linear Rashba SOC, both $B_{x}$ and $B_{z}$ are equally necessary in realizing SDE). By further application of $B_{z}$ the diode efficiency $\eta$ is readily enhanced upto $\sim 21\%$ as demonstrated in Fig. 7(b).
We emphasize that the presence of highly asymmetric behavior in the supercurrent, given a specific set of system parameters, does not necessarily guarantee a corresponding high level of non-reciprocity in the super-currents. As evident from Fig. 6(a), the system might exhibit nearly negligible diode effect ($\eta\sim 0$) even though the parameters are set such that it supports asymmetry in supercurrent.
IV Summary and conclusions
This article analyzes how the interplay between linear and cubic Rashba SOC terms, combined with external Zeeman fields, can optimize the efficiency of SDE in a proximitized one-dimensional nanowire device. Using self-consistent mean-field analysis, we demonstrate that SDE, characterized by nonreciprocal critical currents ($|J_{c}^{+}|\neq|J_{c}^{-}|$), arises from Cooper pair momentum asymmetry in the helical superconducting ground state of this time-reversal and inversion symmetry broken system. Remarkably, we achieve a diode efficiency $\eta\gtrsim 45\%$ with only linear Rashba SOC significantly higher than the previously reported efficiency of $\eta\sim 2\%$ from non-self-consistent analyses Legg et al. (2022). In this case with only linear SOC, at least two magnetic fields are required to obtain a finite diode efficiency: one along the SOC direction and one perpendicular to it. This requirement stems from the fact that linear SOC effects can be eliminated through a spin-dependent gauge transformation Braunecker et al. (2010), resulting in a conventional BCS superconducting state with zero Cooper pair momentum when only one Zeeman field is present. However, the presence of two orthogonal Zeeman fields creates a spiral magnetic field texture via the gauge transformation, leading to momentum-dependent band asymmetry and asymmetric FFLO Cooper pairs Braunecker et al. (2010). In contrast, higher-order SOC effects cannot be eliminated via gauge transformations, allowing for finite diode efficiency with only one Zeeman field. However, including higher-order SOC doesn’t improve the maximum diode efficiency (only $\eta\sim 20\%$), this value remains significant compared to previously reported diode efficiencies Legg et al. (2022); Nadeem et al. (2023). Our results underscore the importance of self-consistently determining the order parameter of the superconducting ground state to fully understand the capability and optimize Rashba nanowire devices as superconducting diodes.
Our findings can be readily experimentally tested in spin orbit coupled Rashba nanowires (Alicea, 2012; Liang and Gao, 2012; Manchon et al., 2015; Bercioux and Lucignano, 2015) such as zinc-blende InSb and wurtzite InAs, with lengths of approximately 100 nm, where cubic Rashba SOC interactions are significant depending on the growth direction Campos et al. (2018). Superconductivity can be proximity induced in these nanowires by common bulk $s$-wave superconductors like Nb or Al. Future research directions include investigating SDE in proximity-induced multichannel Rashba nanowires with overlapping channels Potter and Lee (2010); Lutchyn et al. (2011); Law and Lee (2011); Alicea (2012), extending beyond the single-channel case examined here. Additionally, the influence of topological superconductivity realized in the nanowire and the nanowire device geometry on SDE warrants further exploration Liu et al. (2023); Kutlin and Mel’nikov (2020); Kopasov et al. (2021); Legg et al. (2023); Nesterov et al. (2016); Spånslätt (2018). An intriguing prospect is the potential replication of magnetic field effects using electric fields via the Rashba-Edelstein effect Vignale and Tokatly (2016); Smirnov and Golub (2017); Piatti (2021); Nadeem et al. (2023), which could lead to the realization of field-free SDE.
V Acknowledgments
SKG acknowledges financial support from SERB, Government of India via the Startup Research Grant: SRG/2023/000934 and IIT Kanpur via the Initiation Grant (IITK/PHY/2022116). A.K.N. acknowledges the financial support from Department of Atomic Energy (DAE), Govt. of India, through the project Basic Research in Physical and Multidisciplinary Sciences via RIN4001. The authors acknowledge discussions with Koushik Mandal, Krishnendu Sengupta and Amit Agarwal. DS and SKG utilized the Andromeda server at IIT Kanpur for numerical calculations. SB and AS acknowledge SAMKHYA: High-Performance Computing Facility provided by Institute of Physics, Bhubaneswar, for numerical computations.